### v1

# Extreme resonant eccentricity excitation of stars around merging black-hole binary

Abstract
We study the dynamics of a star orbiting a merging black-hole binary (BHB) in a coplanar triple configuration. During the BHB’s orbital decay, the system can be driven across the apsidal precession resonance, where the apsidal precession rate of the stellar orbit matches that of the inner BHB. As a result, the system gets captured into a state of resonance advection until the merger of the BHB, leading to an extreme eccentricity growth of the stellar orbit. This resonance advection occurs when the inner binary has a non-zero eccentricity and unequal masses. The resonant driving of the stellar eccentricity can significantly alter the hardening rate of the inner BHB, and produce observational signatures to uncover the presence of nearby merging or merged BHBs.
Introduction.— About 90 double compact object mergers have been detected by the LIGO-Virgo-KAGRA Collaboration in recent years LIGO-2021 . Stellar-mass black-hole binary (BHB) mergers can arise through diverse formation channels, including the standard isolated binary evolution Lipunov-1997 ; Lipunov-2007 ; Podsiadlowski-2003 ; Belczynski-2010 ; Belczynski-2016 ; Dominik-2012 ; Dominik-2013 ; Dominik-2015 ; Alejandro-2017 , chemically homogeneous evolution Mandel-2016 ; Marchant-2016 ; duBuisson ; Riley , and multiple-body evolution in the gas disks of active galactic nuclei Baruteau-2011 ; McKernan-2012 ; McKernan-2018 ; Bartos-2017 ; Stone-2017 ; Leigh-2018 ; Secunda-2019 ; Yang-2019 ; Grobner-2020 ; Ishibashi-2020 ; Tagawa-2020 ; Liyaping-2021 ; Ford-2021 ; Samsing-Nature ; Lirixin-2022 ; Lijiaru-2022 . Additionally, various dynamical channels involve strong gravitational scatterings in dense clusters Zwart(2000) ; OLeary(2006) ; Miller(2009) ; Banerjee(2010) ; Downing(2010) ; Ziosi(2014) ; Rodriguez(2015) ; Samsing(2017) ; Samsing(2018) ; Rodriguez(2018) ; Gondan(2018) , tertiary-induced mergers via von Zeipel-Lidov-Kozai (ZLK) oscillations vonZeipel ; Lidov ; Kozai ; Smadar ; Miller-2002 ; Wen-2003 ; Antonini-2012 ; Antonini(2017) ; Silsbee(2017) ; Petrovich-2017 ; Liu-ApJ ; Xianyu-2018 ; Hoang-2018 ; Liu-Quadruple ; Fragione-Quadruple ; Fragione-nulearcluster ; Zevin-2019 ; Liu-HierarchicalMerger or flyby-induced mergers Michaely-2019 ; Michaely-2020 .
In the standard ZLK-induced merger scenario, the tertiary companion must be in a sufficiently inclined orbit to induce eccentricity excitation in the inner binary. However, an alternative eccentricity-pumping mechanism, known as apsidal procession resonance, has been proposed for triples with small mutual inclinations (with no ZLK oscillations) Liu-Yuan ; Liu-2020-PRD . In this mechanism, the apsidal precession rates of both the inner and outer orbits are in a 1:1 commensurability i.e.,
$\dot{\varpi}_{\mathrm{in}}\simeq\dot{\varpi}_{\mathrm{out}},$ (1)
allowing efficient eccentricity transfer between the two orbits. This resonance, driven by the lowest-order post-Newtonian (PN) effect for the inner binary ($\dot{\varpi}_{\mathrm{in}}$) and the Newtonian interaction between the inner and outer orbits ($\dot{\varpi}_{\mathrm{out}}$), can produce merger events with either low-mass or supermassive black-hole tertiary companions Liu-Yuan ; Liu-2020-PRD .
BHB progenitors, potentially numerous yet undetected, could form an abundant population in stellar multiples (triples or quadruples) Tokovinin ; Raghavan ; Fuhrmann ; Moe-2017 . Observationally, triple systems with massive stars are suggested to be less inclined or nearly coplanar, especially if the tertiary is not so faraway Ransom ; Thompson ; Eisner . If the tertiary companion is a visible object (e.g., a star or a pulsar), its motion can aid in the search for inspiral BHBs Suto-1 ; Suto-2 ; Suto-3 ; Liu-Daniel . In this letter, we focus on the secular dynamics of a star orbiting a merging BHB in a coplanar setup. Beginning with an eccentric BHB and a nearly circular outer stellar orbit, we demonstrate that the system can enter an apsidal resonance and be captured into a ‘resonance advection’ state during the inner BHB’s gravitational radiation driven orbital decay. Once captured, the eccentricity of the star grows continually, approaching unity in magnitude.
Orbital evolution.— Consider an inner BHB with masses $m_{1}$, $m_{2}$, and a tertiary companion ($m_{3}$) orbiting the center of mass of the inner bodies. The reduced mass for the inner binary is $\mu_{\mathrm{in}}\equiv m_{1}m_{2}/m_{12}$ where $m_{12}\equiv m_{1}+m_{2}$. Similarly, the outer binary has $\mu_{\mathrm{out}}\equiv(m_{12}m_{3})/(m_{12}+m_{3})$. The semi-major axes and eccentricities are denoted by $a_{\mathrm{in}}$, $a_{\mathrm{out}}$ and $e_{\mathrm{in}}$, $e_{\mathrm{out}}$, respectively. Therefore, the orbital angular momenta of the two orbits are given by $L_{\mathrm{in}}=\mu_{\mathrm{in}}\sqrt{Gm_{12}a_{\mathrm{in}}(1-e_{\mathrm{in}}^{2})}$ and $L_{\mathrm{out}}=\mu_{\mathrm{out}}\sqrt{Gm_{123}a_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})}$.
Figure 1: The evolution of a merging BHB with a test-mass tertiary in a coplanar orbital configuration. The masses of BHB are $m_{1}=46.15M_{\odot}$ and $m_{2}=13.85M_{\odot}$. The initial semimajor axes and eccentricities of the inner and outer binaries are $a_{\mathrm{in},0}=0.4\mathrm{AU}$, $e_{\mathrm{in},0}=0.8$, $a_{\mathrm{out}}=5\mathrm{AU}$ and $e_{\mathrm{out},0}=0.1$, and the longitudes of the periapsis of the two orbits are set to be $\varpi_{\mathrm{in},0}=\varpi_{\mathrm{out},0}=0$. The black (inner binary) and orange (outer binary) lines represent the numerical solutions of the DA secular equations. Key moments during the inner BHB evolution are highlighted in the top two panels (gray dots) with the orbital parameter values labeled. The bottom panel displays the apsidal precession rates given by Eqs. (2)-(3).
To understand the dynamics of the outer binary influenced by the merging BHB, we first consider the case where $m_{3}$ is a test-mass, and thus the back-reaction from the tertiary on the inner binary can be ignored.
The top two panels of Fig. 1 show an example of the orbital evolution of a coplanar triple. The system satisfies the dynamical stability criterion Holman . The results are obtained by integrating the double-averaged (DA; averaging over both the inner and outer orbital periods) secular equations of motion Liuetal-2015 . All Newtonian effects (up to the octupole order), the leading order PN corrections, and the dissipation due to gravitational wave (GW) are included in the calculation. We see that the BHB undergoes orbital decay and eccentricity circularization due to GW radiation. The merger time of the BHB is approximately given by $T_{\mathrm{merger}}=[5c^{5}a_{\mathrm{in},0}^{4}(1-e_{\mathrm{in},0}^{2})^{7/2}]/[256G^{3}m_{12}^{2}\mu_{\mathrm{in}}]$ Peters-1964 .
The bottom panel describes the time evolution of the apsidal precession rates, in which $\dot{\varpi}_{\mathrm{in}}$ is given by the 1st-order PN effect in the inner BHB
$\dot{\varpi}_{\mathrm{in}}=\frac{3G^{3/2}m_{12}^{3/2}}{c^{2}a_{\mathrm{in}}^{5/2}(1-e_{\mathrm{in}}^{2})},$ (2)
and $\dot{\varpi}_{\mathrm{out}}$ is due to the secular perturbation of the inner binary on the outer orbit Liuetal-2015 ; Liu-SMBHB
$\dot{\varpi}_{\mathrm{out}}=\frac{3}{4}\bigg{(}\frac{G\mu_{\mathrm{in}}m_{3}a_{\mathrm{in}}^{2}}{a_{\mathrm{out}}^{3}}\bigg{)}\frac{1}{L_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})^{3/2}}.$ (3)
We see that starting from a quasi-circular orbit, the outer eccentricity can be excited once $\dot{\varpi}_{\mathrm{out}}$ catches up to $\dot{\varpi}_{\mathrm{in}}$. Importantly, rather than a single resonance passage, both frequencies evolve in lockstep until the inner BHB merges, resulting in a significant growth in $e_{\mathrm{out}}$. It’s worth noting that the final $e_{\mathrm{out}}$ becomes quite large, and the DA approximation may be challenged. We employ alternative single-averaged equations of motion to assess its validity in Appendix A.
Figure 2: Phase-space portraits in terms of the canonical variables ($1-\sqrt{1-e_{\mathrm{out}}^{2}}$ and $\Delta\varpi$) at different stages of the orbital evolution for the system in our fiducial model (refer to Fig. 1). In each panel, $\gamma\equiv\dot{\varpi}_{\mathrm{in}}/\dot{\varpi}_{\mathrm{out}}|_{e_{\mathrm{out}}=0}$ and $\dot{\varpi}_{\mathrm{in}}/\dot{\varpi}_{\mathrm{out}}$ are evaluated by Eqs. (13) and (2)-(3) at the time labeled in Fig. 1. The solid gray lines are the contours of constant $\hat{\mathcal{H}}$ (see Eq. Extreme resonant eccentricity excitation of stars around merging black-hole binary). The existence of a separatrix (black line) divides phase-space into circulation zones (zone I and III) and libration zone (zone II). The orange curve represents the numerical solution of the outer binary’s eccentricity vector $\bm{e}_{\mathrm{out}}$, originating at the red dot and ending at the blue dot. To better visualize both evolution trajectory and separatrix, zoom-in subfigures are included in the top-left panels.
Theory.— To understand the nonlinear resonance capture and advection phenomenon shown in Fig. 1, we consider the secular Hamiltonian of the system:
$\mathcal{H}=-\frac{Gm_{1}m_{2}}{2a_{\mathrm{in}}}-\frac{Gm_{12}m_{3}}{2a_{\mathrm{out}}}+\Phi_{\mathrm{N}}+\Phi_{\mathrm{PN}}.$ (4)
Here $\Phi_{\mathrm{N}}$ is the Newtonian potential energy between the inner and outer orbits up to the octupole order, given by
$\Phi_{\mathrm{N}}=\Phi\frac{-2-3e_{\mathrm{in}}^{2}+\frac{15}{8}\varepsilon_{\mathrm{oct}}e_{\mathrm{in}}(4+3e_{\mathrm{in}}^{2})\cos\Delta\varpi}{6(1-e_{\mathrm{out}}^{2})^{3/2}},$ (5)
where
$\Phi=\frac{3}{4}\frac{G\mu_{\mathrm{in}}m_{3}a_{\mathrm{in}}^{2}}{a_{\mathrm{out}}^{3}},$ (6)
and $\varepsilon_{\mathrm{oct}}=[(m_{1}-m_{2})a_{\mathrm{in}}e_{\mathrm{out}}]/[m_{12}a_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})]$ and $\Delta\varpi\equiv\varpi_{\mathrm{out}}-\varpi_{\mathrm{in}}$ with $\varpi_{\mathrm{in}}$, $\varpi_{\mathrm{out}}$ being the longitudes of pericenters. The term $\Phi_{\mathrm{PN}}$ is the PN potential that induces the apsidal precession of the inner binary:
$\Phi_{\mathrm{PN}}=-\frac{3G^{2}m_{1}m_{2}m_{12}}{c^{2}a_{\mathrm{in}}^{2}\sqrt{1-e_{\mathrm{in}}^{2}}}.$ (7)
To examine the dynamics of the outer eccentricity vector $\bm{e}_{\mathrm{out}}$ in a frame corotating with the inner eccentricity vector $\bm{e}_{\mathrm{in}}$, we can rewrite the Hamiltonian as (after dropping the non-essential constant terms)
$\mathcal{H}_{\mathrm{rot}}=\Phi_{\mathrm{N}}-\dot{\varpi}_{\mathrm{in}}L_{\mathrm{out}}.$ (8)
By introducing $\Phi_{0}=\Phi|_{a_{\mathrm{in}}=a_{\mathrm{in},0}}$ with $a_{\mathrm{in},0}$ being the initial semimajor axis, Eq. (8) can be made dimensionless as
$\overline{\mathcal{H}}$ $=\frac{\mathcal{H}_{\mathrm{rot}}}{\Phi_{0}}=A\hat{\mathcal{H}}$
$=A\bigg{[}\frac{B+C\frac{e_{\mathrm{out}}}{1-e_{\mathrm{out}}^{2}}\cos\Delta\varpi}{6(1-e_{\mathrm{out}}^{2})^{3/2}}-\gamma\sqrt{1-e_{\mathrm{out}}^{2}}\bigg{]},$
where the coefficients are defined as
$A\equiv\bigg{(}\frac{a_{\mathrm{in}}}{a_{\mathrm{in},0}}\bigg{)}^{2},$ (10)
$B\equiv-2-3e_{\mathrm{in}}^{2},$ (11)
$C\equiv\frac{15}{8}\frac{m_{1}-m_{2}}{m_{1}+m_{2}}\frac{a_{\mathrm{in}}}{a_{\mathrm{out}}}e_{\mathrm{in}}(4+3e_{\mathrm{in}}^{2}),$ (12)
$\gamma\equiv\frac{\dot{\varpi}_{\mathrm{in}}}{\dot{\varpi}^{\prime}_{\mathrm{out}}}\simeq\frac{4Gm_{12}^{3}}{c^{2}m_{1}m_{2}a_{\mathrm{in}}}\bigg{(}\frac{a_{\mathrm{out}}}{a_{\mathrm{in}}}\bigg{)}^{7/2}\frac{1}{1-e_{\mathrm{in}}^{2}}$ (13)
with $\dot{\varpi}^{\prime}_{\mathrm{out}}=\dot{\varpi}_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})^{2}$. These coefficients evolve slowly as the BHB undergoes orbital decay driven by GW emission.
Fig. 2 shows the dynamical behavior of the vector $\bm{e}_{\mathrm{out}}$ in the phase-space at different stages of the BHB’s orbital decay (marked by the value of $\gamma$). Here, Poincaré variables are adopted, with $(1-\sqrt{1-e_{\mathrm{out}}^{2}})$ and $\Delta\varpi$ representing the canonical momentum and angle, respectively. For each value of $\gamma$, the evolution of $\bm{e}_{\mathrm{out}}$ follows a level curve of constant $\hat{\mathcal{H}}$ (gray line), either circulating freely from 0 to $2\pi$ or librating around a fixed $\Delta\varpi$ (i.e., resonance). The circulating and librating trajectories of $\hat{\mathcal{H}}$ divide the phase-space into different zones, bounded by the separatrix (thick black line) around a given $\gamma$, the trajectory of $\bm{e}_{\mathrm{out}}$ (orange line) can only evolve in one of the zones.
In the beginning, $\gamma<1$ and $\bm{e}_{\mathrm{out}}$ circulates (top left panel of Fig. 2). As the inner BHB decays, $\gamma$ increases. Around $\gamma\simeq 1$, the $\bm{e}_{\mathrm{out}}$ trajectory (orange line) crosses the separatrix, and the circulating trajectory transitions into a librating one, indicating that the system is captured into the resonance (the second panel of Fig. 2). Although the total energy of the system is not conserved, the evolution of $\bm{e}_{\mathrm{out}}$ can be described by a series of ‘frozen’ Hamiltonians for each individual ($a_{\mathrm{in}}$, $e_{\mathrm{in}}$). As $\gamma$ evolves in time, the separatrix gradually shifts to higher values of $e_{\mathrm{out}}$. The librating trajectory is carried along with $\Delta\varpi\simeq 0$ within zone II (termed ‘advection’), accompanied by significant increase in $e_{\mathrm{out}}$. During this ‘adiabatic’ evolution within the separatrix, the phase-space area of the trajectory,
[…]
is conserved. Near the merger of the BHB, the area of zone II shrinks until the separatrix disappears, and the resonant trajectory transitions to circulating eventually (see lower panels).
Note that apsidal precession resonance and the possibility of exciting eccentricity at resonance haven been well recognized in the planetary dynamics literatures, often in the linear eccentricity regime (e.g., Heppenheimer ; Ward ; Lin ; XuWenrui ; Diego ; Suyubo ). However, the distinct advection effect, sweeping through the full range of canonical momentum, is a unique feature of our present study. Three requirements for the system to end up captured into resonance advection have to be satisfied Quillen ; Henrard , including:
i) Adiabatic evolution. The separatrix crossing must be slow. Namely, the timescale of $e_{\mathrm{out}}-$excitation due to resonance should be shorter than the merger time of BHB Liu-Daniel , i.e., $\dot{\varpi}_{\mathrm{out}}T_{\mathrm{merger}}\gg 1$, which is easily satisfied.
ii) Specific direction of time-varying $\gamma$. The resonance encounter must be from $\dot{\varpi}_{\mathrm{out}}>\dot{\varpi}_{\mathrm{in}}$ towards $\dot{\varpi}_{\mathrm{out}}<\dot{\varpi}_{\mathrm{in}}$. For the opposite case, only a one-time jump in the eccentricity due to resonance crossing is excepted Liu-2020-PRD .
iii) Sufficiently small initial eccentricity. The initial eccentricity of the tertiary must be small (e.g., $e_{\mathrm{out},0}\lesssim 0.6$ in our case), so that it can be captured into the resonance while zone III has not yet appeared (see below and Appendix C).
Fig. 3 illustrates the excitation of stellar eccentricity for a range of $(e_{\mathrm{in},0},e_{\mathrm{out},0})$. We find that the cases with $e_{\mathrm{in},0}\gtrsim e_{\mathrm{out},0}$ always give rise to resonant driving of $e_{\mathrm{out}}$, with weak dependence on the initial eccentricities. This behavior can be understood as follows. In the absence of separatrix crossing, as noted before, the phase-space area $\mathcal{A}$ (Eq.14) is an adiabatic invariant. At the time of separatrix crossing, $\mathcal{A}$ is not conserved, and the trajectory can transition into nearby zones Suyubo ; Henrard . For a system initially set with a quasi-circular outer orbit, the trajectory of $\bm{e}_{\mathrm{out}}$ circulates near the bottom part of the phase-space portrait (see the top left panel of Fig. 2). Because there are only two zones at separatrix crossing, the system is guaranteed to be captured into zone II (as in the top panels of Fig. 2). When $0.3\lesssim e_{\mathrm{out}}\lesssim 0.6$, zone III emerges in the transition, and both zone II and III become available. The probabilities of transition to zone $\rm II$ and $\rm III$ are given by
$P_{\rm I\rightarrow\rm II}=-\frac{\partial\mathcal{A}_{\rm II}/\partial\gamma}{\partial\mathcal{A}_{\rm I}/\partial\gamma}=-\frac{\partial\mathcal{A}_{\rm II}}{\partial\mathcal{A}_{\rm I}},$ (15)
$P_{\rm I\rightarrow\rm III}=-\frac{\partial\mathcal{A}_{\rm III}/\partial\gamma}{\partial\mathcal{A}_{\rm I}/\partial\gamma}=-\frac{\partial\mathcal{A}_{\rm III}}{\partial\mathcal{A}_{\rm I}},$ (16)
where $\mathcal{A}_{\rm I(\rm II,III)}$ is the area of the individual zone. Note that $d\mathcal{A}_{\rm I}+d\mathcal{A}_{\rm II}+d\mathcal{A}_{\rm III}=0$. In our case, both $P_{\rm I\rightarrow\rm II}$ and $P_{\rm I\rightarrow\rm III}$ are positive due to the expansion of zone II and zone III and the contraction of zone I at the separatrix crossing. Whether the system transitions to zone II or zone III depends on the initial $\Delta\varpi_{0}$, which affects the initial phase-space area for a given $e_{\mathrm{out},0}$ (see more details in the Appendix B). If $e_{\mathrm{out},0}$ is even higher ($\gtrsim 0.6$), with zone II starting to shrink, capture into zone II becomes impossible as $P_{\rm I\rightarrow\rm II}$ is negative.
Figure 3: The evolution of the outer binary eccentricity as a function of the inner binary semimajor axis. We consider different combinations of ($e_{\mathrm{in},0}$, $e_{\mathrm{out},0}$) in our fiducial model (with the parameters the same as in Fig. 1). All systems satisfy the stability criterion.
As mentioned earlier, a necessary condition for the resonance capture is $\gamma\lesssim 1$ for the initial system. This translate into the constraint that the initial $a_{\mathrm{out}}$ should be larger than a few times of the instability limit Holman (see Fig. 4).
Finite $m_{3}$: modified evolution of merging BHBs.— We now consider the ‘BHB$+$star’ system where the tertiary companion has a finite mass $m_{3}=1M_{\odot}$. In this case, the perturbation from the tertiary companion on the inner binary cannot be ignored. Since the outer binary is allowed to ‘gain’ some eccentricity from the inner BHB at resonance, the inner binary can experience an eccentricity decrease, therefore reducing the orbital decay rate of the BHB. When the initial $e_{\mathrm{out}}$ is large, the inner binary can ‘gain’ some eccentricity, therefore increasing the decay rate.
To explore how the evolution of the BHB can be altered, we begin with the BHB with $m_{1}=46.15M_{\odot}$, $m_{2}=13.85M_{\odot}$, $a_{\mathrm{in},0}=0.4\mathrm{AU}$ and $e_{\mathrm{in},0}=0.8$ in our fiducial model. Based on Fig. 4, we consider the tertiary star with $a_{\mathrm{out}}$ in the range of $(5,10)\mathrm{AU}$ and all possible eccentricities within the stability limit. Random initial $\Delta\varpi_{0}$ values are chosen. The results are presented in panel (A) of Fig. 5. We see that the BHB may experience either slower or faster hardening compared to the isolated binary. The BHB with a longer merger time experiences extra eccentricity decrease due to the tertiary companion, associated with the eccentricity excitation of the tertiary. On the other hand, the BHB with the shorter merger time gains eccentricity from the outer companion, accompanied by decrease in $e_{\mathrm{out}}$.
Figure 4: Parameter space in the $a_{\mathrm{in}}$-$a_{\mathrm{out}}$ plane showing which system may undergo resonance capture and advection during the orbital decay of the inner BHB. We consider different types of BHBs as labeled. Assuming the tertiary body is a test mass with a circular orbit, the cross-hatched region indicates dynamically unstable systems, and the gray-shaded region signifies systems with $\gamma\gtrsim 1$, or $\dot{\varpi}_{\mathrm{in}}\gtrsim\dot{\varpi}_{\mathrm{out}}$ (see Eq. 13). The vertical lines refer to the merger time $T_{\mathrm{merger}}=10^{10}$yrs.
The middle and lower panels of Fig. 5 depict four examples of the eccentricity evolution for different $e_{\mathrm{out},0}$ and $\Delta\varpi_{0}$ values, corresponding to different initial phase-space trajectories of the outer binary. For $e_{\mathrm{out},0}=0.1$ (middle panel), we see that $e_{\mathrm{out}}$ experiences resonant excitation and eventually freezes at a finite value; at the same time, $e_{\mathrm{in}}$ experiences ‘extra’ decrease during the resonance advection leading to a decelerated merger event. For $e_{\mathrm{out},0}=0.4$ (lower panel), the eccentricity evolution is sensitive to the initial value of $\Delta\varpi_{0}$ (or the initial phase-space trajectory); the merger undergoes either advection-deceleration or acceleration, as a result of the probabilistic transition during the separatrix crossing.
Figure 5: Top panel: the evolution of merging BHBs with and without the perturbation of a tertiary star. We use our fiducial model (see Fig. 1 for the parameters) with various $a_{\mathrm{out}}$, $e_{\mathrm{out},0}$ (constrained by Fig. 4) and $\Delta\varpi_{0}$ values. Lower panels: the evolution of orbital eccentricities for the system that have undergone resonance passage or capture. We choose $a_{\mathrm{out}}=5.3\mathrm{AU}$ and $e_{\mathrm{out},0}=0.1$ for case (A) and $a_{\mathrm{out}}=5.4\mathrm{AU}$ and $e_{\mathrm{out},0}=0.4$ for case (B). In each panel, the solid and dashed lines correspond to systems initialized with $\Delta\varpi_{0}=0$ and $\pi$, respectively.
Summary and discussion.— We have shown that a star orbiting around an inspiraling BHB can experience a distinct resonance capture and advection phenomenon, leading to a significant increase in the eccentricity. This resonance occurs when the apsidal precession rates of the inner and outer orbits are comparable. The resonance advection gives rise to the apsidal alignment, enabling a continuous exchange of angular momentum between the two orbits until the BHB merges. The eccentricity growth requires the inner binary to have an unequal masses and an elliptical orbit, and the outer orbit to have a relatively small eccentricity (see Appendix). For systems with initially large outer eccentricities, the resonance capture occurs probabilistically or may not occur at all. In this study, we have restricted to coplanar configurations. Our preliminary study shows that as long as the mutual inclinations between the inner and outer orbits is small ($\lesssim 30^{\circ}$), similar resonance phenomenon can happen.
The dramatic increase in the outer binary eccentricity $e_{\mathrm{out}}$ has a number of applications. The effective exchange of angular momentum between inner and outer binaries alters the evolution of $e_{\mathrm{in}}$, thereby influencing the hardening rate of the BHBs. Additionally, the notable increase in $e_{\mathrm{out}}$ serves as a distinctive signature for revealing hidden BHBs Liu-Daniel that could be observed by upcoming GW detectors like LISA LISA , TianQin TianQin , Taiji TaiJi and B-DECIGO DECIGO . In some cases, the significant $e_{\mathrm{out}}-$excitation could also result in a sufficiently small pericenter distance for the outer stellar orbit, leading to possibly micro tidal disruption events MTDE .
The occurrence rate of the ‘BHB$+$star’ systems studied in this letter is unknown. The progenitor stars of these BHBs often expand to hundreds or thousands of solar radii during their evolution, potentially interacting dynamically with the stellar tertiary. However, low-metallicity massive stellar binaries could remain compact Mandel-2016 ; Marchant-2016 ; duBuisson ; Riley , allowing for dynamically-stable compact triples Alejandro2021 . Alternatively, a BHB might initially form and subsequently capture a long-lived low-mass star.
The resonant eccentricity excitation mechanism studied here can be applied to other astrophysical systems, including a planet orbiting a binary star and a stellar-mass object around a supermassive BHB with or without a gaseous disc. We leave these topics to the future investigations.
Acknowledgement.— B.L. thanks Yubo Su for useful discussions. B.L. gratefully acknowledges support from the European Union’s Horizon 2021 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101065374.
Appendix A A: Equations of motion
Figure 6: Sample orbital evolution of a BHB with a tertiary companion. System parameters are the same as our fiducial example from Fig. 1 in the main text, except for the mass of the tertiary companion (as labeled). The solid and dashed lines denote two integration methods. In the SA calculations, the initial true anomaly of the outer binary is set to be $\pi$.
The secular interaction timescale of the outer orbit driven by the inner binary is $|\dot{\varpi}_{\mathrm{out}}|^{-1}$ (Eq. 3), where
$\dot{\varpi}_{\mathrm{out}}\sim\frac{\mu_{\mathrm{in}}}{m_{12}}\bigg{(}\frac{a_{\mathrm{in}}}{a_{\mathrm{out}}}\bigg{)}^{2}(1-e_{\mathrm{out}}^{2})^{-2}n_{\mathrm{out}}.$ (17)
On the other hand, the pericenter frequency of the outer orbit is
$n_{p}\sim n_{\mathrm{out}}(1-e_{\mathrm{out}}^{2})^{-3/2}.$ (18)
The double-averaged secular equations are valid when $n_{p}\gg\dot{\varpi}_{\mathrm{out}}$, or $\frac{\mu_{\mathrm{in}}}{m_{12}}\big{(}\frac{a_{\mathrm{in}}}{a_{\mathrm{out}}}\big{)}^{2}(1-e_{\mathrm{out}}^{2})^{-1/2}\ll 1$.
To calibrate our different approaches, Fig. 6 presents the examples of the orbital evolution of a BHB with a coplanar companion, obtained by using double-averaged (DA) and single-averaged (SA) secular equations Liuetal-2015 ; Liu-ApJ . The outcomes obtained from these two integration methods exhibit near-identical results. The evolution achieved by the SA equations permits exchanges in both energy and angular momentum. We find weak oscillations in the evolution of $a_{\mathrm{out}}$ in the upper panel when $e_{\mathrm{out}}$ attains a sufficiently large value (depicted by the dashed line). These oscillations can be attributed to the stability of the system.
Appendix B B: Resonance capture and advection
Figure 7: Similar to Fig. 2 in the main text, but we set the initial eccentricity to $e_{\mathrm{out},0}=0.6$ with $\Delta\varpi_{0}=0$ (top two panels) and $\Delta\varpi_{0}=\pi$ (bottom two panels).
Figure 8: The final $e_{\mathrm{out}}$ as a function of $\Delta\varpi_{0}\equiv\varpi_{\mathrm{out},0}-\varpi_{\mathrm{in},0}$ for different values of $e_{\mathrm{out},0}$. The system studied here is from our fiducial model, in which the outer binary is set to $a_{\mathrm{out}}=5\mathrm{AU}$ with the initial eccentricity being specified as labeled.
In this section, we explore the dynamics of the outer binary influenced by the merging inner BHB, initially set with non-zero $e_{\mathrm{out}}$. We again focus on our fiducial model with $m_{1}=46.15M_{\odot}$, $m_{2}=13.85M_{\odot}$, $a_{\mathrm{in},0}=0.4\mathrm{AU}$ and $e_{\mathrm{in},0}=0.8$, and consider the tertiary companion is a test-mass.
Similar to Fig. 2 in the main text, Fig. 7 depicts the dynamic evolution of the vector $\bm{e}_{\mathrm{out}}$ with an initial value of $e_{\mathrm{out},0}=0.6$. As there is no back-reaction from the tertiary on the inner binary, the structure of the phase-space portrait remains the same as Fig. 2, featuring the same contours of constant $\hat{\mathcal{H}}$ and the separatrix for each $\gamma$. However, in contrast to the case with $e_{\mathrm{out},0}\ll 1$, the trajectory of $\bm{e}_{\mathrm{out}}$ here encounters the separatrix around $\gamma\sim 3$ for which zone III already exists. With zone II expanding, both the transition probabilities into zone II and zone III ($P_{\rm I\rightarrow\rm II}$ and $P_{\rm I\rightarrow\rm III}$) are positive (Eqs. 15-16). We find that system initialized with different $\Delta\varpi_{0}$ can undergo either resonance capture or passage, as shown in the upper and lower panels of Fig. 7. The dependence arises because, for a given $e_{\mathrm{out},0}$, the enclosed phase-space areas of $\mathcal{A}_{\rm I(\rm II,III)}$ are different for different $\Delta\varpi_{0}$. Note that for the case of $\Delta\varpi_{0}=0$, the librating evolution trajectory transitions to circulating at $\gamma\simeq 35.1$. This is because zone II starts shrinking in area at this moment, while the phase-space area covered by the evolution trajectory is still conserved. The separatrix crossing becomes unavoidable, and the resonant eccentricity excitation terminates prematurely.
Fig. 8 examines the impact of $e_{\mathrm{out},0}$ and $\Delta\varpi_{0}$ for the resonance passage and advection. For each $e_{\mathrm{out},0}$, the final value of outer binary eccentricity is computed as a function of $\Delta\varpi_{0}$. We see that for the systems with $e_{\mathrm{out},0}\ll 1$, the resonance advection is guaranteed. When $e_{\mathrm{out},0}\gtrsim 0.4$, the distribution of $e_{\mathrm{out},\mathrm{f}}$ imply that resonance advection occurs probabilistically.
Appendix C C: Apsidal resonance: theoretical analysis
As mentioned in the main text, the dimensionless secular Hamiltonian of a coplanar triple is given by
$\hat{\mathcal{H}}=$ $\frac{1}{6(1-e_{\mathrm{out}}^{2})^{3/2}}\bigg{[}B+C\frac{e_{\mathrm{out}}}{1-e_{\mathrm{out}}^{2}}\cos\Delta\varpi\bigg{]}$
$-\gamma\sqrt{1-e_{\mathrm{out}}^{2}},$
where the coefficients are defined as
$B=-|B|=-(2+3e_{\mathrm{in}}^{2}),$ (20)
$C=\frac{15}{8}\frac{m_{1}-m_{2}}{m_{1}+m_{2}}\frac{a_{\mathrm{in}}}{a_{\mathrm{out}}}e_{\mathrm{in}}(4+3e_{\mathrm{in}}^{2}),$ (21)
$\gamma=\frac{4Gm_{12}^{3}}{c^{2}m_{1}m_{2}a_{\mathrm{in}}}\bigg{(}\frac{a_{\mathrm{out}}}{a_{\mathrm{in}}}\bigg{)}^{7/2}\frac{1}{1-e_{\mathrm{in}}^{2}}.$ (22)
By assuming $e_{\mathrm{out}}^{2}\ll 1$, $\hat{\mathcal{H}}$ (Eq. C) can be further simplified to (dropping a non-essential constant and a multiplicity factor)
^$\hat{\mathcal{H}}=-De_{\mathrm{out}}^{2}+Ce_{\mathrm{out}}\cos\Delta\varpi$ (23)
with
$D=\frac{3}{2}|B|-3\gamma.$ (24)
Figure 9: Phase-space portraits in terms of $e_{\mathrm{out}}$ and $\Delta\varpi$. Three lines are given by Eq. (25) when ^$\hat{\mathcal{H}}>0$ (blue), ^$\hat{\mathcal{H}}=0$ (black) and ^$\hat{\mathcal{H}}<0$ (red). Here, $\beta$ is evaluated by the initial orbital parameters shown in Fig. 1 in the main text.
During the orbital decay of the inner BHB, the coefficients $|B|$ (of order a few) and $C$ (typically $\ll 1$) decrease in time, while $\gamma$ starts from $\lesssim 1$ and increases to large value. For the system studied in this letter, $D>0$ is valid at the early stage of the orbital decay, the evolution of the outer binary eccentricity is thus determined by
$e_{\mathrm{out}}=\frac{1}{2}\bigg{[}\beta\cos\Delta\varpi\pm\sqrt{\beta^{2}\cos^{2}\Delta\varpi+4\hat{\mathcal{H}}^{\prime}}\bigg{]},$ (25)
where $\beta=C/D$ and ^=^$\hat{\mathcal{H}}^{\prime}=\hat{\mathcal{H}}/D$.
Fig. 9 shows the solutions of Eq. (25). Similar to Fig. 2 in the main text, we see that a separatrix (black line; corresponding to ^$\hat{\mathcal{H}}=0$) devides the phase-space into two zones, above which the outer binary eccentricity follows the circulating trajectory (blue line; as ^$\hat{\mathcal{H}}>0$) and below which $e_{\mathrm{out}}$ is librating (red line; as ^$\hat{\mathcal{H}}<0$). Importantly, the peak value of $e_{\mathrm{out}}$ of the separatrix is given by $e_{\mathrm{out},\mathrm{peak}}=\beta$, which occurs at $\Delta\varpi=0$.
Therefore, to have secure resonance capture, the requirements can be summarized as:
i) Initially, $D$ should be greater than zero, which translates to $\gamma\lesssim 1$;
ii) Adiabatic evolution.
iii) Specific direction of time-varying $\gamma$.
iv) Sufficiently small initial eccentricity. Based on the theoretical analysis above, the requirement for $e_{\mathrm{out},0}$ can now be well quantified as
$e_{\mathrm{out},0}$ $\lesssim e_{\mathrm{out},\mathrm{peak}}~{}(\mathrm{at~{}~{}\gamma\sim 1})$
$\simeq\frac{15}{8}\frac{m_{1}-m_{2}}{m_{1}+m_{2}}\frac{a_{\mathrm{in}}}{a_{\mathrm{out}}}e_{\mathrm{in}}(4+3e_{\mathrm{in}}^{2})~{}(\mathrm{at~{}~{}\gamma\sim 1}).$